OpenStudy (anonymous):

Solve for x. 3(x + 5) - 2(x + 2) = 20 1 9 13

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Callisto

OpenStudy (candyme):

3(x+5)-2(x+2)=20 First thing you want is to multiply out those Parentheses

OpenStudy (candyme):

Use the distributive property

OpenStudy (candyme):

3(x)+(3)(5)+(-2)(x)+(-2)(2)=20 3x+15-2x-4=20 Now you want to combine like terms 3x-2x=x and 15-4=11 x+11=20 Now subtract the 11 from both sides x=20-11 x=9

OpenStudy (anonymous):

thank you